### v1

# Learning General Representation of 12-Lead Electrocardiogram With a Joint-Embedding Predictive Architecture

Abstract
We propose a self-supervised learning method for 12-lead Electrocardiogram (ECG) analysis, named ECG Joint Embedding Predictive Architecture (ECG-JEPA). ECG-JEPA employs a masking strategy to learn semantic representations of ECG data. Unlike existing methods, ECG-JEPA predicts at the hidden representation level rather than reconstructing raw data. This approach offers several advantages in the ECG domain: (1) it avoids producing unnecessary details, such as noise, which is common in standard ECG; and (2) it addresses the limitations of naïve L2 loss between raw signals. Another key contribution is the introduction of a special masked attention tailored for 12-lead ECG data, Cross-Pattern Attention (CroPA). CroPA enables the model to effectively capture inter-patch relationships. Additionally, ECG-JEPA is highly scalable, allowing efficient training on large datasets. Our code is openly available https://github.com/sehunfromdaegu/ECG_JEPA.
1 Introduction
Electrocardiograms (ECG) provide a non-invasive method to measure the electrical activity of the heart over time, serving as a crucial tool for diagnosing various cardiac conditions. While numerous supervised learning models have been developed to detect heart diseases using ECG data [1, 2, 3], these models often face significant performance degradation when applied to data distributions different from those on which they were trained.
Self-supervised learning (SSL) has emerged as a powerful paradigm for learning general representations across various domains, including natural language processing [4, 5, 6], computer vision [7, 8, 9], and video analysis [10, 11]. Despite its promise, applying SSL to ECG data presents unique challenges. For instance, data augmentation, which is essential in many SSL architectures, is more complex for ECG than for computer vision (CV) data. Simple transformations like rotation, scaling, and flipping, effective in CV, can distort the physiological meaning of ECG signals. Additionally, ECG recordings often contain artifacts and noise, which cause autoencoder-based SSL models to struggle with reconstructing raw signals. These architectures may also miss visually subtle but diagnostically critical features, such as P-waves and T-waves, which are imperative for diagnosing certain cardiac conditions.
In this work, we propose a novel ECG Joint-Embedding Predictive Architecture (ECG-JEPA) tailored for 12-lead ECG data, effectively addressing the aforementioned challenges. ECG-JEPA utilizes a transformer architecture to capture the semantic meanings of ECG. By masking several patches of the ECG data, ECG-JEPA predicts the abstract representations of the missing patches, indicating a high-level understanding of the data. Additionally, we develop a novel masked-attention for multi-lead ECG data, coined Cross-Pattern Attention (CroPA). CroPA enables the transformer architecture to effectively capture the relationships between patches in multi-lead ECG.
In this work, we demonstrate that:
  • ECG-JEPA significantly improves linear evaluation and fine-tuning on classification tasks compared to existing SSL methods without hand-crafted augmentations.
  • CroPA enhances the model’s ability to capture inter-lead relationships, leading to better performance on downstream tasks.
  • ECG-JEPA can also recover important ECG features, including heart rate and QRS duration, which are classical indicators used in ECG evaluation. This is the first work to demonstrate that learned representations can effectively recover various ECG features.
  • ECG-JEPA is highly scalable, allowing efficient training on large datasets. For instance, ECG-JEPA is trained for only 100 epochs, yet outperforms other ECG SSL models on most downstream tasks, taking approximately 26 hours on a single RTX 3090 GPU.
Figure 1: 12-lead ECG. The standard 12-lead ECG captures electrical activity from different angles around the heart, providing comprehensive cardiac assessment. Artifacts and noise, such as powerline interference and baseline wander, are common in ECG recordings. The figure shows an example of a 12-lead ECG with baseline wander.
2 Background
Self-Supervised Learning (SSL) facilitates learning abstract representations from input data without the need for labeled data, which is particularly beneficial in medical domains where labeled data is scarce and expensive. SSL leverages inherent data patterns to learn useful representations, allowing models to adapt to various downstream tasks with greater robustness to data imbalances [12].
2.1 Generative Architectures
Generative architectures involve reconstructing an input $x$ from its degraded version $x^{\prime}$ using an encoder-decoder framework. The idea behind generative architectures is that the ability to reconstruct the clean data from a corrupted one indicates a good understanding of the data. The encoder maps the perturbed input $x^{\prime}$ into a latent representation, which the decoder then uses to generate the original input $x$ [13].
2.2 Joint-Embedding Predictive Architectures
Joint-Embedding Predictive Architectures (JEPA) [14] process pairs $x$ and their corrupted versions $x^{\prime}$ to obtain representations $z$ and $z^{\prime}$ through encoders. Unlike generative architectures that predict in the input space, JEPA predicts in the hidden representation space by reconstructing $z$ from $z^{\prime}$. This approach effectively avoids the challenge of predicting unpredictable details, which is common in biological signals.
2.3 Electrocardiogram (ECG)
The electrocardiogram (ECG) is a non-invasive diagnostic method that records the heart’s electrical activity over time using electrodes placed on the skin. The standard 12-lead ECG captures electrical activity from multiple angles. These 12 leads are categorized into limb leads (I, II, III), augmented limb leads (aVR, aVL, aVF), and chest leads (V1-V6). Each lead provides unique information about the heart’s electrical activity, offering a comprehensive view that aids in diagnosing various cardiac conditions. Refer to Figure 1 for an illustration.
In this work, we use only 8 leads (I, II, V1-V6) as the remaining 4 leads (III, aVR, aVL, aVF) can be derived from linear combinations of the 8 leads. This choice maintains the necessary diagnostic information while optimizing computational efficiency.
Figure 2: Key ECG Features.
2.4 ECG Features
[…]
3 Methodology
ECG-JEPA is trained by predicting masked representations of ECG data within the hidden representation space, using only a partial view of the input. The proposed architecture utilizes a student-teacher network framework, as illustrated in Figure 3. We subdivide the multi-channel ECG into non-overlapping patches and sample a subset of these patches for masking. However, reconstructing the raw signals of masked patches can be particularly challenging in the ECG domain due to the prevalence of noise in biological signals. Instead, our model predicts the masked patches in the hidden representation space, where this challenge can be effectively addressed. We validate the quality of the learned representations through various downstream tasks, including linear probing, fine-tuning on classification tasks, and ECG feature extraction tasks.
3.1 Patch Masking
Let $x\in\mathbb{R}^{L\times T}$ represent a multi-lead ECG of length $T$ with $L$ channels. We divide the interval $[0,T]$ into $N$ non-overlapping subintervals of length $t$. Each subinterval in each channel constitutes a patch of $x$, resulting in $L\times N$ patches. The masking strategy in multi-lead ECG must be carefully chosen because patches in different leads at the same temporal position are highly correlated, potentially making the prediction task too easy. To address this, we mask all patches across different leads in the same temporal space. With this in mind, we employ two masking strategies: random masking and multi-block masking.
In random masking, we randomly select a percentage of subintervals to mask, while in multi-block masking, we select multiple consecutive subintervals to mask. Note that we allow these consecutive subintervals to overlap, which requires the model to predict much longer sequences of representations. In this paper, we use both masking strategies to evaluate the effectiveness of ECG-JEPA, with a random masking ratio of $(0.6,0.7)$ and a multi-block masking ratio of $(0.175,0.225)$ with a frequency of 4. The unmasked patches serve as the contextual input for the student networks, while the masked patches are the ones for which we aim to predict the representations.
The patches are then converted into sequences of token vectors using a linear layer, and augmented with positional embeddings. For positional encoding, we employ the conventional 2-dimensional sinusoidal positional embeddings for the student and teacher networks, while we use 1-dimensional sinusoidal positional embeddings for the predictor network.
Figure 3: ECG-JEPA. The student network processes only non-masked patches of time series, whereas the teacher network processes the entire time series. Then the predictor predicts the representation of masked patches processed by the teacher. Note that both student and teacher network process the multi-channel time series, while predictor gets only a single-channel time series. In this picture, there are $L=3$ leads with $N=5$ subintervals and $Q=3$ unmasked subintervals.
3.2 Teacher, Student, and Predictor
ECG-JEPA consists of three main components: the teacher network, the student network, and the predictor network. Both the teacher and student networks are based on standard transformer architectures. The weights of the teacher network are updated through an exponential moving average (EMA) of the student network, with details provided in the supplementary materials. The predictor network, a smaller transformer, operates on single-channel representations.
The teacher network handles the entire $L\times N$ patches, generating fully contextualized $L\times N$ representations. The student network, however, processes only $L\times Q$ visible (unmasked) patches, where $Q<N$ represents the number of visible time intervals. These $L\times Q$ representations from the student are then concatenated with the (learnable) mask tokens, resulting in $L\times N$ representations. Subsequently, each lead’s representations are passed to the predictor, which processes single-channel representations. The predictor’s output, the predicted representations of the target patches, is compared with the target representations using a smooth L1 loss function.
3.3 Cross-Pattern Attention (CroPA)
Biological signals often require simultaneous observation across multiple channels. In a 12-lead ECG, abnormalities must be consistent across multiple leads to diagnose specific cardiac pathologies. Therefore, the meaning of a single patch should be understood in the context of other patches within the same channel and across different channels. In contrast, patches in different channels and temporal spaces are often neglected when interpreting a 12-lead ECG.
To address this, we introduce Cross-Pattern Attention (CroPA), a masked self-attention tailored for multi-lead ECG data. CroPA allows each patch to attend only to patches within the same channel and temporal space (Figure 4). This method enhances our model’s effectiveness compared to the standard attention mechanism, where patches attend to all other patches indiscriminately.
4 Related Work
This section provides an overview of existing SSL methods that have been applied to ECG data. We compare ECG-JEPA with these methods on various downstream tasks in the next section.
4.1 Contrastive Learning
Contrastive learning has been widely explored in various research works for its effectiveness in unsupervised representation learning. Contrastive Predictive Coding (CPC) [15] aims to capture useful data representations by predicting future observations in the latent space. It uses sequential data where the “true future" is treated as the positive sample, and other observations within the same sequence are treated as negative samples. SimCLR [7] and MoCo [16] also leverage contrastive approaches, with positive samples being augmented versions of the same image and negative samples being other images. SimCLR relies on large batch sizes and extensive data augmentation, while MoCo uses a momentum encoder to maintain a dynamic dictionary of encoded samples, enabling more stable training with smaller batch sizes.
4.2 Generative Architectures
BERT [4] pioneered generative architectures in natural language processing (NLP), inspiring similar methods in other domains. Masked Autoencoders (MAE) [8] have been developed for computer vision, where a portion of the input image is masked, and the model is trained to reconstruct these missing parts. Recently, [17] pointed out that generative architectures often prioritize learning the principal subspaces of the data, which may limit their effectiveness in capturing semantic representations for perceptual tasks.
Figure 4: Cross-Pattern Attention (CroPA). The patch in the middle attends only to the colored patches, which are in the same lead and temporal space.
4.3 Joint-Embedding Predictive Architectures
Data2vec [18] introduced an SSL architecture that predicts the average of the encoder’s final layers’ outputs, which was further improved in efficiency in data2vec 2.0 [19]. Notable models like I-JEPA [9] and V-JEPA [11] predict the hidden representations of masked patches, enabling semantic representation learning for images and videos, respectively.
4.4 Self-Supervised Learning for ECG
Several attempts have been made to train SSL models for 12-lead ECG data. Contrastive Multi-segment Coding (CMSC) [20] is a contrastive method that divides an ECG into two segments, encouraging close representations for compatible segments while pushing apart incompatible ones. Contrastive Predictive Coding (CPC) [15] has also been used for ECG data in [21], predicting future representations of the ECG. A significant disadvantage of CPC is that it is not scalable to large datasets, as the LSTM modules inside CPC make the training process extremely slow. Recently, [22] utilized masked autoencoders for ECG data, proposing temporal and channel masking strategies, coined as Masked Time Autoencoder (MTAE) and Masked Lead Autoencoder (MLAE), respectively. Similarly, [23] proposed ST-MEM, an MAE for ECG data that masks random time intervals for each lead. One possible drawback of MLAE and ST-MEM is that predicting some patches can be relatively easy considering that 12 leads signals are highly correlated.
5 Experimental Settings
This section provides an overview of the experimental settings. Details of all experiments can be found in the supplementary materials.
5.1 Pretraining Datasets
Training SSL models with large datasets is crucial for developing generalized representations. However, most previous works have used relatively small datasets, with the exception of [23], where an SSL model was trained with a large number of 12-lead ECGs. Following [23], we use the Chapman [24], Ningbo [25], and CODE-15 [26] datasets for pretraining ECG-JEPA. The Chapman and Ningbo datasets collectively consist of 45,152 10-second 12-lead ECGs at 500Hz. CODE-15 includes 345,779 12-lead ECGs from 233,770 patients at 400Hz, with 143,328 being 10-second recordings. After excluding recordings with missing values, we have 43,240 ECGs from Chapman and Ningbo and 130,900 ECGs from CODE-15. All ECGs were resampled to 250Hz for consistency.
5.2 Downstream Datasets
We use the PTB-XL [27] and CPSC2018 [28] datasets to evaluate the performance of ECG-JEPA on downstream tasks. PTB-XL contains 21,837 clinical 10-second 12-lead ECG records from 18,885 patients, recorded at 500Hz and annotated with 71 diagnostic labels, which are aggregated into five superclasses. We use these superclass labels for our experiments. The CPSC2018 dataset includes 6,877 12-lead ECG recordings with nine annotated cardiac conditions. These datasets are multi-label in nature, where each recording can have multiple labels simultaneously. The details of the datasets are provided in the supplementary materials.
5.3 Pretraining
After resampling the ECG data to 250Hz, each 10-second multi-lead ECG has $T=2500$ time points. We divide the interval $[0,T]$ into $N=50$ non-overlapping subintervals of length $t=50$. The model is trained for 100 epochs without any data augmentations, and the final checkpoint is used for downstream tasks.
5.4 Architecture
We use vision transformer (ViT-Base) architectures [29] for the student and teacher network, and a smaller standard transformer architecture for the predictor network. The student and teacher network process the multi-lead ECG, while the predictor processes each lead independently to reconstruct the masked representations. See supplementary materials for detailed architecture specifications.
5.5 Downstream Tasks
We conduct extensive experiments to show that ECG-JEPA effectively captures semantic representations. We evaluate its performance on classification tasks through linear probing and fine-tuning. Additionally, we evaluate the performance on the low-shot setting, and on the reduced-lead setting where the downstream dataset is restricted to have a single or two leads.
Next, we validate the expressive power of the learned representations by predicting ECG features, including heart rate and QRS duration. To the best of our knowledge, this is the first work showing that the learned representations can recover various ECG features. The successful prediction of these features provides compelling evidence that the representations are not only informative but also capture essential and clinically relevant characteristics, making them highly reliable for ECG analysis.
In ECG analysis, it is common for a single recording to have multiple labels simultaneously, making datasets like PTB-XL and CPSC2018 multi-label tasks. However, many previous studies have converted this into a multi-class classification problem by restricting the dataset to subsets with only a single label. To ensure a fair comparison, we will pretrain the methods using publicly available code and evaluate their performance on the multi-label classification task. If the code is not available, we will convert our task into a multi-class problem and compare the results with the reported performance in the literature.
6 Experiments
In this section, we assess the performance of the learned representations on various downstream tasks. Our experiments aim to demonstrate that the learned representations are generalizable and effective in capturing essential ECG features. We compare ECG-JEPA with several state-of-the-art SSL methods.
For classification tasks, we use AUC (Area Under the ROC Curve) and F1 scores as our primary evaluation metrics. AUC evaluates model performance across all possible thresholds, providing a comprehensive and stable measure of discriminative ability. In contrast, the F1 score balances precision and recall at a fixed threshold. In our experiments, we use the AUC as our main metric, as it provides a more robust evaluation of the model’s performance across varying classification thresholds, ensuring that the results are not overly dependent on a single decision boundary.
For multi-label classification, the AUC is computed by averaging the scores of binary classifications for each label. For multi-class classification, the AUC is calculated using the average of one-vs-rest approach. In both cases, the F1 score is computed using a macro-averaged approach that averages the F1 scores of all classes.
In most cases, ECG-JEPA outperforms other SSL methods that leverage hand-crafted augmentations, demonstrating its effectiveness in learning general representations of ECG data. In experiments, E-JEPArb and E-JEPAmb refer to ECG-JEPA models trained with random masking and multi-block masking strategies, respectively.
6.1 Linear Evaluation
Tables 1 and 2 present the results of our linear evaluation on the PTB-XL and CPSC2018 datasets. We train a linear classifier on top of the frozen representations for 10 epochs, and evaluate its performance on downstream tasks. As shown in the tables, ECG-JEPA consistently outperforms other SSL methods, demonstrating superior efficiency and effectiveness with substantially reduced computational resources.
Table 1: Linear evaluation on multi-label task.
PTB-XL CPSC2018
Method Epochs AUC F1 AUC F1
ST-MEM 800 0.896 0.662 0.964 0.752
SimCLR 300 0.866 0.624 0.890 0.523
CMSC 300 0.802 0.472 0.767 0.206
CPC 100 0.620 0.167 0.687 0.091
E-JEPArb 100 0.906 0.690 0.969 0.769
E-JEPAmb 100 0.912 0.712 0.971 0.789
Table 2: Linear evaluation on multi-class task.
PTB-XL CPSC2018
Method Epochs AUC F1 AUC F1
MoCo v3111Score as reported in [23]. 800 0.739 0.142 0.712 0.080
MTAE [^1] Score as reported in [23]. 800 0.807 0.437 0.818 0.349
MLAE [^1] Score as reported in [23]. 800 0.779 0.382 0.794 0.263
ST-MEM 800 0.888 0.566 0.973 0.805
SimCLR 300 0.842 0.496 0.918 0.624
CMSC 300 0.796 0.442 0.787 0.391
CPC 100 0.600 0.201 0.672 0.210
E-JEPArb 100 0.894 0.616 0.974 0.805
E-JEPAmb 100 0.896 0.628 0.973 0.819
6.2 Fine-tuning
Fine-tuning is another method to evaluate the quality of learned representations, as it tests the model’s ability to adapt its pre-trained features to new tasks. In the fine-tuning process, we add a linear classification head at the end of the encoder and train the whole network for 10 epochs. Fine-tuning can potentially enhance performance beyond what is achieved with linear evaluation alone.
Table 3 presents the results of fine-tuning on the PTB-XL and CPSC2018 datasets. ECG-JEPA is compared with other SSL methods in a multi-class classification setting. The results indicate that ECG-JEPA achieves the highest AUC and F1 scores on PTB-XL and the highest AUC on CPSC2018. These results highlight ECG-JEPA’s effectiveness in fine-tuning scenarios, demonstrating its ability to leverage transfer learning by refining and adapting its learned representations for improved performance in specific classification tasks.
Table 3: Fine-tuning on multi-class task.
PTB-XL CPSC2018
Method Epochs AUC F1 AUC F1
MoCo v3111Score as reported in [23]. 800 0.913 0.644 0.967 0.838
MTAE [^1] Score as reported in [23]. 800 0.910 0.613 0.961 0.769
MLAE [^1] Score as reported in [23]. 800 0.915 0.625 0.973 0.816
CMSC [^1] Score as reported in [23]. 800 0.877 0.510 0.938 0.717
ST-MEM 800 0.929 0.668 0.977 0.820
SimCLR 300 0.905 0.650 0.934 0.693
CPC [^2] We did not fine-tune CPC due to its slow training process. 100 - - - -
E-JEPArb 100 0.944 0.710 0.980 0.821
E-JEPAmb 100 0.937 0.680 0.983 0.799
6.3 Low-shot Linear Evaluation
Table 4 presents the performance comparison on the low-shot task. Low-shot learning is particularly challenging, as models must generalize effectively with limited labeled data. Given the difficulty and resource-intensive nature of obtaining labeled data in medical research, low-shot learning represents a realistic and critical scenario in the medical field. In this experiment, we evaluate the performance of ECG-SSL models on the PTB-XL multi-label task with only 1% and 10% of the training set, while keeping the test set fixed. As shown in the table, ECG-JEPA demonstrates a clear advantage over other SSL methods, with its effectiveness becoming particularly evident in low-shot learning tasks. This suggests that ECG-JEPA can be particularly well-suited for transfer learning where labeled data is scarce.
Table 4: Low-shot linear evaluation on the multi-label PTB-XL task. The mean and standard deviation of macro AUC are reported for 1% (192 samples) and 10% (1923 samples) of the training set, selected three times independently.
PTB-XL
Method Epochs 1% 10%
ST-MEM 800 0.807 $\pm$ 0.005 0.872 $\pm$ 0.001
SimCLR 300 0.803 $\pm$ 0.002 0.843 $\pm$ 0.001
CMSC 300 0.750 $\pm$ 0.008 0.792 $\pm$ 0.001
CPC 100 0.523 ±$\pm 0.006$ 0.560 $\pm$ 0.005
E-JEPArb 100 0.836 $\pm$ 0.006 0.887 $\pm$ 0.000
E-JEPAmb 100 0.843 $\pm$ 0.004 0.894 $\pm$ 0.003
6.4 Reduced Lead Evaluation
Because transformer architectures can accept variable input lengths, we can evaluate ECG-JEPA’s performance on reduced lead tasks. In this experiment, we perform a linear evaluation on the PTB-XL multi-label task using reduced leads. Specifically, we utilize only a single lead (Lead II) and two leads (Lead II and V1), training linear classifiers on top of the learned representations for 10 epochs. We then compare ECG-JEPA’s performance with ST-MEM, another model supporting reduced lead evaluation. Table 5 presents the results of this reduced lead linear evaluation. Notably, ECG-JEPA performs well even in the reduced lead setting. This robustness is particularly valuable since most daily mobile devices typically output only one or two leads, highlighting ECG-JEPA’s potential for practical applications in mobile health monitoring.
Table 5: Reduced lead evaluation. Linear evaluation on PTB-XL multi-label task with reduced lead settings.
1-Lead 2-Lead
Method AUC F1 AUC F1
ST-MEM 0.832 0.571 0.848 0.597
E-JEPArb 0.846 0.596 0.877 0.647
E-JEPAmb 0.849 0.593 0.880 0.657
6.5 ECG Feature Extraction
ECG features are essential for diagnosing and monitoring cardiac conditions. We evaluate the model’s ability to extract key ECG features, including heart rate and QRS duration, from the learned representations of the PTB-XL dataset. Unlike classification tasks, which are highly perceptual, ECG features are directly related to the morphology of the signal. For example, heart rate can be measured by the average distance between R-peaks, and QRS duration can be measured by the average width of the QRS complex.
Several methods have been proposed to segment ECG signals [30, 31, 32, 33]. Segmentation models can be used to extract ECG features. In this experiment, we employed a publicly available segmentation model [33] to extract ECG features from the PTB-XL dataset, which served as our ground truth labels. We then trained a linear regression model on the learned representations to predict these ECG features, using the mean squared error (MSE) as the loss function.
Table 6 presents the performance comparison, showing the means and standard deviations of the absolute differences between the extracted and predicted values across all samples in the PTB-XL test set.
Interestingly, although the model’s learned representations are implied to capture high-level features based on their strong performance in linear evaluation tasks across various datasets, they still retain the ability to recover low-level information, such as ECG features. This dual capability of capturing both high-level and low-level information is noteworthy and highlights the model’s versatility.
Table 6: ECG feature extraction. Means and standard deviations of absolute differences between the predicted and extracted values for heart rate and QRS duration. Small values indicate better performance. The final row provides the overall mean and standard deviation for both features across all samples in the PTB-XL test set.
Method Heart Rate (BPM) QRS Dur. (ms)
ST-MEM 1.35 $\pm$ 2.38 4.60 $\pm$ 4.16
SimCLR 1.87 $\pm$ 2.81 6.14 $\pm$ 5.80
CMSC 7.20 $\pm$ 7.43 10.12 $\pm$ 9.98
CPC 11.40 $\pm$ 11.04 11.55 $\pm$ 11.55
E-JEPArb 1.54 $\pm$ 2.62 4.81 $\pm$ 4.29
E-JEPAmb 1.45 $\pm$ 2.44 4.41 $\pm$ 4.08
Mean 75.01 $\pm$ 17.65 90.48 $\pm$ 17.02
7 Ablation Study
7.1 Effect of CroPA
Table 7 presents the results of our evaluation of the effectiveness of CroPA. CroPA introduces a “human-like" inductive bias, enabling the model to be trained more efficiently on multi-lead ECG data. Without CroPA, models may require more epochs to converge. For a fair comparison, we trained ECG-JEPA with and without CroPA for 100 and 200 epochs and compared their performance on the PTB-XL multi-class task. The results show that CroPA considerably improves the model’s performance, demonstrating its effectiveness in capturing inter-lead relationships and enhancing the model’s ability to learn meaningful representations.
Table 7: Effect of CroPA. Linear evaluation (lin) and fine-tuning (ft) results on PTB-XL multi-class task with ECG-JEPArb and ECG-JEPAmb. In both cases, CroPA improves the model’s performance.
lin ft
Mask CroPA Epochs AUC AUC
Random x 100 0.888 0.930
Random x 200 0.887 0.927
Random o 100 0.894 0.944
Multi-block x 100 0.872 0.924
Multi-block x 200 0.886 0.914
Multi-block o 100 0.896 0.937
7.2 Masking Ratio
Table 8 presents the performance of ECG-JEPA in linear evaluation with different masking ratios and strategies. The results indicate that the model benefits from a high masking ratio. Notably, multi-block masking is advantageous for linear evaluation, while random masking is more effective for fine-tuning, as indicated in Table 3. Although random masking with a ratio of (0.7, 0.8) achieves better performance in the PTB-XL multi-label task, a masking ratio of (0.6, 0.7) performs better in other tasks. Therefore, we chose the latter for our main experiments.
Table 8: Effect of masking strategy. Linear evaluation results on PTB-XL multi-label task using different masking ratios and strategies.
Mask Ratio Freq. AUC F1
Random (0.3, 0.4) 1 0.884 0.652
Random (0.4, 0.5) 1 0.904 0.698
Random (0.5, 0.6) 1 0.906 0.697
Random (0.6, 0.7) 1 0.906 0.690
Random (0.7, 0.8) 1 0.909 0.706
Multi-block (0.10, 0.15) 4 0.904 0.678
Multi-block (0.15, 0.20) 4 0.905 0.687
Multi-block (0.175, 0.225) 4 0.912 0.712
8 Future Work
Our method has potential applications in various physiological multivariate signals, such as EEG and EMG. These signals share characteristics with ECG, including their multivariate nature and high dimensionality, making them suitable for our proposed approach.
Another promising direction involves the integration of multi-modal data. For example, multi-lead ECG data could be combined with other diagnostic modalities, such as chest X-rays (CXR), to provide a more comprehensive understanding of a patient’s condition.
One significant challenge in pursuing these extensions is the scarcity of large-scale datasets. Addressing this limitation is crucial for advancing and validating our method across diverse applications.
9 Conclusion
We proposed ECG-JEPA, a novel SSL method tailored for 12-lead ECG data. By utilizing a JEPA coupled with the innovative relative positional encoding method, CroPA, ECG-JEPA effectively learns meaningful representations of ECG signals. This approach addresses the challenges posed by noise and artifacts in ECG data, demonstrating substantial improvements over existing SSL methods in various downstream tasks, with the added benefit of significantly faster convergence.
Our extensive experimental evaluations reveal that ECG-JEPA outperforms state-of-the-art SSL methods across several tasks, including linear evaluation, fine-tuning, low-shot learning, and ECG feature extraction. Moreover, our investigation into the use of 8 leads, as opposed to the full 12-lead ECG, indicates that this reduction does not compromise performance while optimizing computational efficiency. This finding is particularly significant for applications constrained by limited computational resources.
Appendix A Supplementary Materials
A.1 Experimental Details on ECG-JEPA
Hyperparameters for ECG-JEPA pretraining, linear evaluation, and fine-tuning are provided in Tables 9, 10, and 11, respectively. In ECG-JEPAmb, the number of visible patches in ECG-JEPAmb varies more than in ECG-JEPArb, resulting in higher GPU memory usage. Consequently, we reduced the batch size to 64 to fit the model on a single NVIDIA RTX 3090 GPU. Interestingly, ECG-JEPAmb benefits from larger learning rates, even with the halved batch size.
For fine-tuning process, the actual learning rate is calculated as $lr=base\text{\_}lr\times batchsize/256$, following the heuristic by [34].
Table 9: Pretraining Settings for ECG-JEPA.
[…]
optimizer AdamW AdamW
learning rate 2.5e-5 5e-5
weight decay 0.05 0.05
batch size 128 64
learning rate schedule cosine decay cosine decay
warmup epochs 5 5
epochs 100 100
drop path 0.1 0.1
Table 10: Linear Evaluation Settings
config value
optimizer AdamW
learning rate 5e-4
weight decay 0.05
batch size 32
learning rate schedule cosine decay
warmup epochs 3
epochs 10
Table 11: Fine-tuning Settings.
config value
optimizer AdamW
base learning rate 1.0e-4
weight decay 0.05
batch size 16
learning rate schedule cosine decay
warmup epochs 3
epochs 10
A.2 Pretraining Other SSL Models
Besides pretraining ECG-JEPA, we also pretrained other models, including CMSC [20], CPC [15], and SimCLR [7] using the same datasets as ECG-JEPA.
For CMSC and CPC, we adhered to the original architecture and hyperparameters. SimCLR utilized a ResNet50 [35] encoder with an output dimension of 2048. CMSC and SimCLR were pretrained for 300 epochs, selecting the best checkpoint at 100, 200, or 300 epochs based on linear evaluation performance on the PTB-XL multi-label setting. Due to the slow training process, CPC was pretrained for only 100 epochs, taking approximately 9 days on a single NVIDIA RTX 3090 GPU due to the LSTM module in the model. For ST-MEM [23], we employed the publicly available checkpoint pretrained for 800 epochs.
Given SimCLR’s sensitivity to data augmentations, we applied several that work well empirically: baseline shift (adding a constant to all leads), baseline wander (low-frequency noise), Gaussian noise (random noise), powerline noise (50Hz noise), channel resize, random crop, and jump noise (sudden jumps). These augmentations aimed to enhance the robustness of the model to various signal distortions.
A.3 Downstream Datasets Details
Table 12, 13, 14, and 15 show the distribution of the PTB-XL and CPSC2018 datasets in both multi-label and multi-class settings.
The PTB-XL dataset is stratified into ten folds, where the first eight folds are used for training, the ninth fold for validation, and the tenth fold for testing. In our experiments, we used the first nine folds for training and the tenth fold for testing, as we did not observe overfitting during linear evaluation and fine-tuning.
For the CPSC2018 dataset, only the training set is publicly available, which is stratified into seven folds. We used the first six folds for training and the seventh fold for testing, omitting the validation set. The original CPSC2018 dataset consists of 6,877 ECG recordings, but we excluded recordings with a length of less than 10 seconds, resulting in 6,867 ECG recordings.
Table 12: PTB-XL multi-label distribution. The sum of samples in each class exceeds the total number of ECG recordings due to the multi-label nature of the dataset.
# ECG Norm MI STTC CD HYP
Multi-label 21799 9514 5469 5235 4898 2649
Train 19230 8551 4919 4714 4402 2387
Test 2158 963 550 521 496 262
Table 13: PTB-XL multi-class distribution. ECGs with multiple labels are excluded from the PTB-XL multi-label dataset.
# ECG Norm MI STTC CD HYP
Multi-class 16244 9069 2532 2400 1708 535
Train 14594 8157 2276 2158 1524 479
Test 1650 912 256 242 184 56
Table 14: CPSC2018 multi-label distribution. The sum of samples in each class exceeds the total number of ECG recordings due to the multi-label nature of the dataset.
# ECG Norm PVC AF LBBB STE 1AVB PAC STD RBBB
Multi-label 6867 918 1220 235 220 721 614 699 868 1854
Train 5989 805 1059 206 197 632 534 615 742 1616
Test 878 113 161 29 23 89 80 84 126 238
Table 15: CPSC2018 multi-class distribution. The number of ECG recordings for each superclass in the PTB-XL dataset. Note that the sum of samples in each class exceeds the total number of ECG recordings due to multi-label the nature of the dataset.
# ECG Norm PVC AF LBBB STE 1AVB PAC STD RBBB
Multi-class 6391 918 975 178 185 685 531 606 783 1530
Train 5577 805 849 159 169 600 459 534 671 1331
Test 814 113 126 19 16 85 72 72 112 199
A.4 Exponential Moving Average
The teacher network is initialized as a copy of the student network and is updated using an exponential moving average (EMA) of the student’s weights. The EMA is computed as follows:
$\theta_{\text{teacher}}^{i}=\beta_{i}\theta_{\text{teacher}}^{i-1}+(1-\beta_{i})\theta_{\text{student}}^{i}$
where $i$ denotes the current training iteration, and $\beta_{i}$ is a momentum parameter that evolves during training. The momentum parameter $\beta_{i}$ is computed as:
$\beta_{i}=\text{ema}_{0}+\frac{i\cdot(\text{ema}_{1}-\text{ema}_{0})}{\text{iterations\_per\_epoch}\cdot\text{epochs}}$
Here, $\text{ema}_{0}$ and $\text{ema}_{1}$ represent the initial and final values of the momentum parameter, respectively. For our implementation, $\text{ema}_{0}=0.996$ and $\text{ema}_{1}=1.0$.
A.5 Sufficiency of 8 Leads in ECG Analysis
In a standard 12-lead ECG, leads III, aVR, aVL, and aVF can be derived from leads I and II, as shown by the following equations:
[…]
$\text{aVR}=-\frac{(\text{I}+\text{II})}{2}$
$\text{aVL}=\text{I}-\frac{\text{II}}{2}$
$\text{aVF}=\text{II}-\frac{\text{I}}{2}$
These derivations demonstrate that leads III, aVR, aVL, and aVF do not provide additional independent information beyond what is captured by leads I and II. Therefore, using only 8 leads (I, II, V1, V2, V3, V4, V5, and V6) can be sufficient for effective ECG analysis. This simplification not only reduces the computational load but also maintains the diagnostic integrity of the analysis.
A.6 Comparison with 12-Lead Model
We now investigate the practical sufficiency of using 8 leads for ECG-JEPA pretraining. To evaluate the impact of this reduction, we trained models using both 8 leads and 12 leads and compared their performance on the linear evaluation of a multi-label task for PTB-XL.
Table 16 presents the results of this comparison using ECG-JEPArb. As expected, the performance difference between the 8-lead and 12-lead models is minimal, indicating that using 8 leads is sufficient for effective pretraining without significant loss of information.
Table 16: Comparison of 8-Lead and 12-Lead Models.
Model epochs AUC F1
8-Lead 100 0.906 0.690
12-Lead 100 0.905 0.699